A boat which has a speed of 5 km/h in still water crosses a river of width 1 km along the shortest possible path in 15 minutes. The velocity of the river was in km/h is ?????? (ANS-3Km/h)

Grade:12

1 Answers

valay
9 years ago
v w.r.t river (v r)is 5 km/hr  . time taken is ¼ hr . shortest distance is the joining two banks making 0 degree with y axis
let v r make r angle with y axis 
1/v r cosr =1/4
cos r = 4/5
r = 37 deg
for it to move along y axis , x component of v must be zero 
v r sin r = v of river 
5*3/5 = 3= v of river